# DKDS: A Benchmark Dataset of Degraded Kuzushiji Documents with Seals for Detection and Binarization

### v1

@@ 3.1 Raw Data Collection
Category Total Training Testing
Kuzushiji Documents 50 40 10
Seals 128 $100\times 2$ 28
Seals per Document - 4 – 6 2 – 3
Figure 6: Examples of verification and manual correction in the dataset construction workflow: “Raw” refers to the input images, “Initial” denotes the initial binarization produced by Ju et al. [ju2024three], and “Corrected” represents the final ground-truth after verification and manual correction. The trained Kuzushiji expert performed verification to determine which elements should be preserved (e.g., marginal notes, voiced sound symbols, periods) and which should be removed (e.g., stains).
Figure 7: Challenging examples for the text and seal detection task: The first row presents seals suffering from ink fading, while the second row shows seals overlapping with Kuzushiji characters or other seals.
3.2 Construction Workflow
We constructed our dataset by producing color document images and their corresponding binarized ground-truth images from the raw data, and the overall workflow is shown in Fig. 5. To create the final document images, we combined 128 background removed seal images with 50 Kuzushiji document images. Specifically, 40 Kuzushiji documents were selected for training, and the remaining 10 for testing. As detailed in Table 1, the 128 seal images were divided into 100 for the 40 training Kuzushiji documents, and 28 for the 10 test Kuzushiji documents. To increase the cases of overlap between seals and Kuzushiji characters in the training set, we duplicated the 100 training seal images to create 200 seals, which were then randomly added onto the 40 training Kuzushiji documents. As a result, each Kuzushiji document image in the training set contains approximately 4–6 seals, while those in the test set contain about 2–3 seals.
To generate the binarization ground-truth, we employed a combined approach of neural network generation followed by manual verification and correction. As shown in Fig. 5, we first applied the model proposed by Ju et al. [ju2024three] to binarize the raw Kuzushiji document images (without added seals), producing the initial binarization ground-truth. Subsequently, a trained Kuzushiji expert manually verified these initial results. Even in the absence of seal interference, the initial binarization often contained errors. As detailed in Fig. 6, the expert identified issues such as missing marginal notes, voiced sound symbols, periods, and unremoved stains. These errors were then manually corrected to create the final binarization ground-truth. Notably, this verification step was critical and required a trained Kuzushiji expert to ensure the validity of the process.
@@ 4.1 Evaluation Tracks
We define two task tracks in the proposed dataset: (1) text and seal detection, and (2) document binarization. As shown in Fig. 7, seals may overlap with Kuzushiji characters or other seals, and suffer from ink fading, all of which can significantly reduce detection accuracy, making this task highly challenging. In the document binarization task, the goal is to remove seals while preserving, or even restoring, the underlying Kuzushiji characters. This task becomes particularly difficult when the characters and seals overlap, as shown in Fig. 8.
Furthermore, both tasks hold important research significance. Text and seal detection serves as a critical preliminary step for downstream applications, including Kuzushiji character recognition and seal processing, while document binarization aims to improve the accuracy of subsequent OCR systems. By addressing these challenges, the proposed dataset encourages the development of robust models capable of processing real-world scenarios in Kuzushiji document analysis.
Figure 8: Challenging examples for the document binarization task: The left column shows the input Kuzushiji document images, while the right column presents the corresponding binarized images.
@@ 4.2 Evaluation Metrics
For the text and seal detection task, we adopted standard object detection evaluation metrics [lin2014microsoft], including model parameters (Params), floating-point operations (FLOPs), average precision at a 50% intersection over union (IoU) threshold ($\text{AP}_{50}$), and average precision across IoU thresholds from 50% to 95% ($\text{AP}_{50:95}$). Model parameters reflect the model’s size and complexity, FLOPs quantify computational cost, and these average precision (AP) metrics provide a comprehensive measure of detection accuracy.
For the document binarization task, we employed classical evaluation metrics [gatos2009icdar] for quantitative comparison, including the F-measure (FM), pseudo F-measure (p-FM), peak signal-to-noise ratio (PSNR), and distance reciprocal distortion (DRD). To provide a more comprehensive assessment, we also adopted the Average-Score Metric (ASM) proposed by Ju et al. [ju2025efficient], which is calculated as follows:
$ASM=\dfrac{FM+p\text{-}FM+PSNR+(100-DRD)}{4}.$ (1)
A higher ASM value indicates better overall performance, providing a single unified metric for model ranking.
Figure 9: The architecture of the GAN-based method used.
@@ 4.3 Baselines
For the text and seal detection task, we employed the YOLO series of object detection models, including YOLOv8 [jocher2023yolov8], YOLOv9 [wang2024yolov9], YOLOv10 [wang2024yolov10], and YOLO11 [jocher2024yolo11]. As one-stage detectors, these models reduce computational complexity and model parameters while maintaining high detection accuracy, achieving a good balance between inference speed and performance. In addition, they support end-to-end training, provide fast inference, and present strong generalization capabilities, making them widely adopted in object detection tasks.
@@
For the document binarization task, we first applied several traditional binarization algorithms to the Kuzushiji documents, including Niblack [niblack1985introduction], Otsu [otsu1979threshold], and Sauvola [sauvola2000adaptive], which can be directly employed without any training.
@@
Considering that seals in the pre-modern Japanese documents are typically red, while Kuzushiji characters are black and the paper background is pale yellow, we introduced a preprocessing step to remove seal regions based on K-Means clustering [mcqueen1967some] before applying the binarization algorithms. This preprocessing effectively reduces the interference caused by red seals and enhances the quality of subsequent binarization results. Specifically, we designed and implemented three combined approaches: K-Means + Otsu, K-Means + Niblack, and K-Means + Sauvola.
@@
In addition, to provide a deep learning-based baseline, we employed a GAN [goodfellow2020generative]-based model to generate binarized Kuzushiji document images. The network architecture for this model is shown in Fig. 9. Since directly inputting the entire image (approximately $2100\times 3200$ pixels) would require excessive GPU memory, we divided each document image into multiple $512\times 512$ patches; the details of this division are described in Section 5.1.2. These patches were then fed into a U-Net [ronneberger2015u] architecture with an EfficientNet-B5 [tan2019rethinking] backbone, which served as the generator. A simplified PatchGAN [isola2017image, zhu2017unpaired] discriminator was adopted to distinguish between real and fake (generated) images. Furthermore, following Suh et al. [suh2022two] and Ju et al. [ju2024three], we employed a loss function that combines the Wasserstein GAN with Gradient Penalty (WGAN-GP) loss [gulrajani2017improved] and an additional Binary Cross-Entropy (BCE) loss.
Table 2: Quantitative comparison of various YOLO models for text and seal detection on the DKDS dataset.
Model Params FLOPs $\text{AP}_{50}^{\text{Kuzushiji}}$ $\uparrow$ $\text{AP}_{50:95}^{\text{Kuzushiji}}$ $\uparrow$ $\text{AP}_{50}^{\text{Seal}}$ $\uparrow$ $\text{AP}_{50:95}^{\text{Seal}}$ $\uparrow$
YOLOv8m [jocher2023yolov8] 25.86M 79.1G 96.4% 71.2% 99.1% 86.2%
YOLOv9m [wang2024yolov9] 20.16M 77.5G 96.3% 71.7% 97.2% 81.4%
YOLOv10m [wang2024yolov10] 16.49M 64.0G 96.2% 71.4% 99.1% 85.7%
YOLO11m [jocher2024yolo11] 20.05M 68.2G 97.8% 74.1% 98.5% 85.7%
Figure 10: Visual comparison of text and seal detection results predicted by different models with the corresponding ground-truth annotations. The input images are “200003803_00028_1” (bottom) and “200003803_00028_2” (top).
Table 3: Quantitative comparison of various methods for document binarization on the DKDS dataset.
[…]
Niblack [niblack1985introduction] 39.13 41.14 8.44dB 79.70 27.25
Otsu [otsu1979threshold] 63.01 63.31 11.76dB 37.69 50.10
Sauvola [sauvola2000adaptive] 87.87 90.99 18.34dB 7.01 72.55
GAN [goodfellow2020generative] 98.11 98.14 26.53dB 0.82 80.49
Figure 11: Visual comparison of results generated by different methods on Kuzushiji document “200003803_00028_2” from the DKDS test set. The input image, ground-truth, and method outputs are shown for comparison.
Figure 12: Visual comparison of results generated by different methods on Kuzushiji document “200003803_00029_1” from the DKDS test set. The input image, ground-truth, and method outputs are shown for comparison.
@@ 5.1 Implementation Details
To ensure a fair performance comparison on the DKDS dataset, all models were trained and evaluated on the same hardware (i.e., NVIDIA RTX 3090 GPU). All experiments were conducted using the Ubuntu operating system, Python 3.9, and the PyTorch framework.
@@ 5.1.1 Text and Seal Detection
For the text and seal detection task, the YOLO models were trained for 100 epochs with a batch size of 16 and an input resolution of $640\times 640$ pixels. Model training was performed using the SGD optimizer [robbins1951stochastic] with an initial learning rate of 0.01. All YOLO models were initialized with weights pre-trained on the COCO dataset [lin2014microsoft].
@@ 5.1.2 Document Binarization
For the document binarization task, traditional binarization algorithms (i.e., Niblack [niblack1985introduction], Otsu [otsu1979threshold], and Sauvola [sauvola2000adaptive]) first convert the input image into a grayscale image as a preprocessing step. The Niblack algorithm uses a $25\times 25$ sliding window with $k=0.8$, while the Sauvola algorithm employs the same window size. For the K-Means clustering method [mcqueen1967some], the number of clusters is set to $K=3$. After clustering, median filtering with a kernel size of 5 is applied to smooth the results and remove isolated noise points.
For the GAN-based method [goodfellow2020generative], the training data were preprocessed by dividing each original image into $512\times 512$ overlapping patches with a 30% overlap ratio. To further expand the training set, we applied data augmentation, including scaling (factors 0.75, 1.0, 1.25, and 1.5) and rotations (0°and 270°). After preprocessing and augmentation, the 40 original training images were expanded into a total of 71,458 patches of size $512\times 512$ pixels. Both the generator and discriminator were optimized using the Adam [kingma2014adam] optimizer with a learning rate of $2\times 10^{-4}$ and $\beta$ coefficients of $(0.5,0.999)$. The model was trained for 10 epochs with a batch size of 16. At inference, the predicted outputs were binarized using a 0.5 threshold.
@@ 5.2.1 Quantitative Comparison
We trained and evaluated various YOLO series models on our proposed DKDS dataset, using the medium (M) size variant for all models. Quantitative comparison results are summarized in Table 2.
For the text (i.e., Kuzushiji character) detection, all models achieved over 96% in $\text{AP}_{50}$, indicating that they can accurately detect most Kuzushiji characters. Under stricter IoU thresholds (i.e., $\text{AP}_{50:95}$), YOLO11m achieved the highest score of 74.1%, outperforming the other models by about 2–3%. This demonstrates the superior accuracy of YOLO11m in localizing Kuzushiji character bounding boxes.
For the seal detection, both YOLOv8m and YOLOv10m obtained the highest $\text{AP}_{50}$ of 99.1%, indicating near-perfect detection performance under the loose IoU criterion. In addition, YOLOv8m reached the highest $\text{AP}_{50:95}$ of 86.2%, demonstrating strong localization accuracy under stricter IoU thresholds.
In terms of model size, YOLOv10m presents the lowest Params of 16.49M and FLOPs of 64.0G, making it the most lightweight model.
Meanwhile, YOLO11m achieves the best balance between accuracy and efficiency, with 20.05M parameters, 68.2G FLOPs, $\text{AP}_{50}^{Kuzushiji}$ of 97.8%, and $\text{AP}_{50}^{Seal}$ of 98.5%. Therefore, YOLO11m can be considered an effective model for the DKDS dataset.
@@ 5.2.2 Visualization
In the first row, YOLOv10m failed to detect a Kuzushiji character at the bottom of the third row from the left, while the other models detected all characters correctly. However, all models incorrectly predicted stains in the upper-left corner as Kuzushiji characters, which is the main factor affecting model performance. Furthermore, although our ground-truth annotations do not include positional information for marginal notes, both YOLOv8m and YOLO11m successfully detected several marginal notes, as shown on the far right.
In the second row, the ink on the Kuzushiji character in the upper-left corner is fading, which poses a challenge for detection. None of the YOLO models except YOLO11m correctly detected this character, demonstrating YOLO11m’s superior capability in Kuzushiji character detection.
@@ 5.3.1 Quantitative Comparison
We employed traditional binarization algorithms, traditional algorithms combined with K-means clustering, and deep learning-based (i.e., GAN) methods as baselines for the document binarization task. The experimental results are summarized in Table 3.
Overall, the traditional binarization algorithms presented poor performance. Specifically, Niblack’s method [niblack1985introduction] performed the worst, as indicated by its high DRD value, reflecting strong sensitivity to noise in the DKDS dataset. Otsu’s method [otsu1979threshold], a simple global thresholding method, achieved moderate results, while Sauvola’s method [sauvola2000adaptive] significantly outperformed both by leveraging locally adaptive thresholding to process complex backgrounds more effectively.
Furthermore, applying K-means clustering [mcqueen1967some] as a preprocessing step to remove red interference (e.g., seals) further improved performance. For instance, the ASM increased from 27.25 to 28.49, from 50.10 to 69.57, and from 72.55 to 73.09 for K-means + Niblack, K-means + Otsu, and K-means + Sauvola, respectively.
In contrast, the deep learning–based method (i.e., GAN) demonstrated superior performance, achieving the FM of 98.11, p-FM of 98.14, PSNR of 26.53 dB, DRD of 0.82, and an ASM of 80.49. The high FM, p-FM, and low DRD values indicate that the generated binary images are highly readable. However, the relatively lower PSNR suggests that, although the Kuzushiji character outlines are clear, some noise or imperfections remain along the edges.

### v2

@@ 3.2 Construction Workflow
Category Total Training Testing-E Testing-D
Document 50 40 10 10
Seal 128 $100\times 2$ 28 $28\times 5$
S/D - 4 – 6 2 – 3 14
S/D denotes the number of seals per document.
@@ 4.1 Evaluation Tracks
We define two task tracks in the proposed dataset: (1) text and seal detection, and (2) document binarization. As shown in Fig. 7, seals may overlap with Kuzushiji characters or with other seals, and suffer from ink fading. These degradations can lead to incomplete seal detection or significantly reduce Kuzushiji character detection accuracy, making the detection task highly challenging. In the document binarization task, the goal is to remove seals while preserving, or even restoring, the underlying Kuzushiji characters. This task becomes particularly difficult when the characters and seals overlap, as shown in Fig. 8.
Furthermore, both tasks hold important research significance. Text and seal detection serves as a critical preliminary step for downstream applications, including Kuzushiji character recognition and seal processing, while document binarization aims to improve the accuracy of subsequent OCR systems. By addressing these challenges, the proposed DKDS dataset encourages the development of robust models capable of processing real-world scenarios in Kuzushiji document analysis.
@@ 4.2 Evaluation Metrics
For the text and seal detection task (Track 1), we adopt standard evaluation metrics commonly used in text detection [karatzas2015icdar, liao2020real], including the number of model parameters (Params), floating-point operations (FLOPs), precision (P, in %), recall (R, in %), and F-measure (F, in %).
For the document binarization task (Track 2), we employ classical evaluation metrics [gatos2009icdar] for quantitative comparison, including the F-measure (FM), pseudo F-measure (p-FM), peak signal-to-noise ratio (PSNR), and distance reciprocal distortion (DRD). To provide a more comprehensive assessment, we also adopt the Average-Score Metric (ASM) proposed by Ju et al. [ju2025efficient, ju2025mfegan], which is calculated as follows:
$\text{ASM}\!=\!\frac{1}{4}(\text{FM}\!+\!\text{p-FM}\!+\!\text{PSNR}\!+\!(100\!-\!\text{DRD)}),$ (1)
where a higher ASM value indicates better overall performance, providing a single unified metric.
@@ 4.3 Baselines
For the text and seal detection task (Track 1), we employ the YOLO series of object detection models, including YOLOv8 [jocher2023yolov8], YOLOv9 [wang2024yolov9], YOLOv10 [wang2024yolov10], and YOLO11 [jocher2024yolo11]. As one-stage detectors, these models reduce computational complexity and model parameters while maintaining high detection accuracy, achieving a good balance between inference speed and performance. In addition, they support end-to-end training, provide fast inference, and present strong generalization capabilities, making them widely adopted in object detection tasks.
@@
For the document binarization task (Track 2), we apply several traditional binarization algorithms to the Kuzushiji documents, including Niblack [niblack1985introduction], Otsu [otsu1979threshold], and Sauvola [sauvola2000adaptive], which can be directly employed without any training.
@@
Considering that seals in the pre-modern Japanese documents are typically red, while Kuzushiji characters are black and the paper background is pale yellow, we introduce a preprocessing step to remove seal regions based on K-Means clustering [mcqueen1967some] before applying the binarization algorithms. This preprocessing effectively reduces the interference caused by red seals and enhances the quality of subsequent binarization results. Specifically, we implement three combined approaches: K-Means + Otsu, K-Means + Niblack, and K-Means + Sauvola.
To provide deep learning-based baselines, we employ two SOTA GAN-based methods (i.e., Suh et al. [suh2022two] and Ju et al. [ju2024three]) to evaluate the proposed DKDS dataset.
@@
In addition, we employ a cGAN [isola2017image]-based method to generate binarized Kuzushiji document images. The network architecture for this model is shown in Fig. 9. Since directly inputting the entire image (approximately $2100\times 3200$ pixels) would require excessive GPU memory, we divide each document image into multiple $512\times 512$ patches; the details of this division are described in Section 5.1.2. These patches are then fed into a U-Net [ronneberger2015u] architecture with an EfficientNet-B5 [tan2019rethinking] backbone, which serves as the generator. A simplified PatchGAN [isola2017image, zhu2017unpaired] discriminator is adopted to distinguish between real and fake (generated) images. Furthermore, following Suh et al. [suh2022two] and Ju et al. [ju2024three], we employ a loss function that combines the Wasserstein GAN with Gradient Penalty (WGAN-GP) loss [gulrajani2017improved] and an additional Binary Cross-Entropy (BCE) loss.
Figure 7: Challenging examples in text and seal detection. The first row presents seals suffering from ink fading, while the second row shows seals overlapping with Kuzushiji characters or other seals.
Figure 8: Challenging examples in document binarization, where Kuzushiji characters overlap with seals.
Figure 9: The architecture of the cGAN-based method used.
@@ 5.1 Implementation Details
To ensure a fair performance comparison on the DKDS dataset, all models were trained and evaluated on the same hardware (i.e., NVIDIA GeForce RTX 3090 GPU). All experiments were conducted using the Ubuntu operating system, Python 3.9, and the PyTorch framework.
@@ 5.1.1 Text and Seal Detection
For the text and seal detection task (Track 1), the YOLO models were trained for 100 epochs with a batch size of 16 and an input resolution of $640\times 640$ pixels. Model training was performed using the SGD optimizer [robbins1951stochastic] with an initial learning rate of 0.01. All YOLO models were initialized with weights pre-trained on the COCO dataset [lin2014microsoft].
Table 2: Quantitative comparison of medium-sized YOLO models for text and seal detection on the DKDS Testing-E set.
Model Params FLOPs Kuzushiji Seal
P R F P R F
YOLOv8 (Ultralytics) [jocher2023yolov8] 25.86M 79.1G 95.5 89.1 92.2 95.6 96.4 96.0
YOLOv9 (ECCV2024) [wang2024yolov9] 20.16M 77.5G 95.9 89.9 92.8 99.9 96.4 98.1
YOLOv10 (NeurIPS2024) [wang2024yolov10] 16.49M 64.0G 93.5 89.9 91.7 99.6 96.4 98.0
YOLO11 (Ultralytics) [jocher2024yolo11] 20.05M 68.2G 97.1 92.4 94.7 97.0 92.9 94.9
Table 3: Quantitative comparison of medium-sized YOLO models for text and seal detection on the DKDS Testing-D set.
Model Params FLOPs Kuzushiji Seal
P R F P R F
YOLOv8 (Ultralytics) [jocher2023yolov8] 25.86M 79.1G 90.5 87.9 89.2 92.9 81.4 86.8
YOLOv9 (ECCV2024) [wang2024yolov9] 20.16M 77.5G 94.0 87.1 90.4 88.4 81.4 84.8
YOLOv10 (NeurIPS2024) [wang2024yolov10] 16.49M 64.0G 90.4 87.1 88.7 96.4 77.1 85.7
YOLO11 (Ultralytics) [jocher2024yolo11] 20.05M 68.2G 94.7 89.1 91.8 98.5 84.3 90.8
Figure 10: Visual comparison of text and seal detection results predicted by different models, together with the corresponding ground-truth annotations. The input image is “200003803_00028_2” from the DKDS Testing-D set.
@@ 5.1.2 Document Binarization
For the document binarization task (Track 2), traditional binarization algorithms (i.e., Niblack [niblack1985introduction], Otsu [otsu1979threshold], and Sauvola [sauvola2000adaptive]) first convert the input images into grayscale as a preprocessing step. The Niblack algorithm uses a $25\times 25$ sliding window with $k=0.8$, while the Sauvola algorithm employs the same window size. For the K-Means clustering method [mcqueen1967some], the number of clusters is set to $K=3$. After clustering, median filtering with a kernel size of $5\times 5$ is applied to smooth the results and remove isolated noise points.
For the two SOTA GAN-based methods, we follow their original training data preprocessing strategies, resulting in 325,608 patches of size $256\times 256$ pixels and 240 downscaled full-page images of $512\times 512$ pixels. In terms of model architecture, we adopt a U-Net [ronneberger2015u] generator for the method of Suh et al. [suh2022two], and employ a U-Net++ [zhou2019unet++] generator for the method of Ju et al. [ju2024three]. Both methods use EfficientNet-B5 [tan2019rethinking] as the encoder backbone.
For our cGAN [isola2017image]-based method, the training data are preprocessed by dividing each original image into $512\times 512$ overlapping patches with a 30% overlap ratio. To further expand the training set, we apply data augmentation, including scaling (factors 0.75, 1.0, 1.25, and 1.5) and rotations (0°and 270°). After preprocessing and augmentation, the 40 original training images are expanded into a total of 71,458 patches of size $512\times 512$ pixels. Both the generator and discriminator are optimized using the Adam [kingma2014adam] optimizer with a learning rate of $2\times 10^{-4}$ and $\beta$ coefficients of $(0.5,0.999)$. The model is trained for 10 epochs with a batch size of 16. At inference, the predicted outputs are binarized using a threshold of 0.5.
@@ 5.2.1 Quantitative Comparison
We train and evaluate various YOLO series models on the proposed DKDS dataset, using the medium (M) size variant for all models, and the quantitative comparison results are shown in Tables 2 and 3.
On the Testing-E set, all models demonstrate strong overall performance. For Kuzushiji text detection, YOLOv11 achieves the best results with an F-measure of 94.7, maintaining an excellent balance between a precision of 97.1 and a recall of 92.4, and significantly outperforms YOLOv8, YOLOv9, and YOLOv10. In contrast, for seal detection, YOLOv9 achieves the highest performance with an F-measure of 98.1, closely followed by YOLOv10 with an F-measure of 98.0 and YOLOv8 with an F-measure of 96.0, while YOLOv11 exhibits comparatively lower performance with an F-measure of 94.9.
The Testing-D set contains a larger number of seals per document, which markedly increases the overall detection difficulty. For Kuzushiji text detection, YOLOv11 continues to achieve the highest F-measure of 91.8, demonstrating robust text modeling and detection capability even under complex degradation conditions. For seal detection, YOLOv11 shows an even more pronounced advantage, attaining an F-measure of 90.8 and clearly outperforming other models, including YOLOv8 with an F-measure of 86.8, YOLOv9 with 84.8, and YOLOv10 with 85.7.
Compared with Testing-E, the recall of seal detection decreases markedly for all models on Testing-D, indicating that incomplete seal detection is the primary source of the increased difficulty observed in Testing-D.
Regarding model efficiency, YOLOv10 has the smallest parameter count at 16.49M and the lowest computational cost at 64.0G FLOPs. Meanwhile, YOLOv11 demonstrates an excellent balance between efficiency and performance while maintaining high detection accuracy.
@@ 5.2.2 Visualization
It can be seen that YOLOv10m fails to detect a Kuzushiji character located at the bottom of the third column from the left, whereas the other models successfully detect all characters. In addition, all YOLO models incorrectly predicted stains in the upper-left corner as Kuzushiji characters, which constitutes a primary factor affecting detection performance. Furthermore, although the ground-truth annotations do not include positional information for marginal notes, all YOLO models detect several marginal notes to varying extents, as shown on the far right of the figure.
Table 4: Quantitative comparison of various methods for document binarization on the DKDS Testing-E set.
[…]
Otsu [otsu1979threshold] 63.01 63.31 11.76dB 37.69 50.10
Niblack [niblack1985introduction] 39.13 41.14 8.44dB 79.70 27.25
Sauvola [sauvola2000adaptive] 87.87 90.99 18.34dB 7.01 72.55
Suh et al. (PR2022) [suh2022two] 187.30M 93.09 93.09 20.99dB 3.12 76.01
Ju et al. (KBS2024) [ju2024three] 191.46M 95.80 95.81 23.14dB 1.76 78.25
Ours 31.22M 98.11 98.14 26.53dB 0.82 80.49
Table 5: Quantitative comparison of various methods for document binarization on the DKDS Testing-D set.
[…]
Otsu [otsu1979threshold] 60.51 60.80 11.39dB 40.89 47.95
Niblack [niblack1985introduction] 37.52 39.41 8.28dB 82.29 25.73
Sauvola [sauvola2000adaptive] 84.00 87.04 17.08dB 9.60 69.63
K-means [mcqueen1967some] + Otsu [otsu1979threshold] 68.10 69.04 13.37dB 33.32 54.30
K-means [mcqueen1967some] + Niblack [niblack1985introduction] 35.98 37.04 8.27dB 82.66 24.66
K-means [mcqueen1967some] + Sauvola [sauvola2000adaptive] 70.50 69.98 15.80dB 14.69 60.40
Suh et al. (PR2022) [suh2022two] 187.30M 91.95 91.96 20.10dB 3.85 75.04
Ju et al. (KBS2024) [ju2024three] 191.46M 94.86 94.90 22.11dB 2.27 77.40
Ours 31.22M 97.08 97.13 24.58dB 1.38 79.35
Figure 11: Visual comparison of results generated by different methods on Kuzushiji document “200003803_00027_1” from the DKDS Testing-D set. The input image, ground-truth image, and various method outputs are shown for comparison.
@@ 5.3.1 Quantitative Comparison
We employ traditional binarization algorithms, traditional algorithms combined with K-means clustering, and deep learning-based (i.e., GAN) methods as baselines for the document binarization task. The experimental results are summarized in Tables 4 and 5.
On the Testing-E set, the traditional binarization algorithms exhibit generally poor performance. Specifically, Niblack’s method [niblack1985introduction] performs the worst, as indicated by its high DRD value, reflecting strong sensitivity to noise in the DKDS dataset. Otsu’s method [otsu1979threshold], a simple global thresholding method, achieves moderate results, while Sauvola’s method [sauvola2000adaptive] significantly outperforms both by leveraging locally adaptive thresholding to process complex backgrounds more effectively. Furthermore, applying K-means clustering [mcqueen1967some] as a preprocessing step to remove red interference (e.g., seals) further improves performance. For instance, the ASM increases from 27.25 to 28.49, from 50.10 to 69.57, and from 72.55 to 73.09 for K-means + Niblack, K-means + Otsu, and K-means + Sauvola, respectively. These results indicate that when the number of seals is limited, color-based clustering effectively mitigates seal interference for traditional thresholding methods.
However, when traditional methods are applied to the Testing-D set, which contains a larger number of seals, the limitations of combining K-means clustering with thresholding methods become more evident. Although K-means preprocessing still provides some improvement for simple thresholding approaches, overall performance deteriorates significantly. In particular, the FM, p-FM, and DRD metrics for both K-means + Otsu and K-means + Sauvola exhibit noticeable degradation. This suggests that in more complex scenarios, clustering-based binarization strategies struggle to robustly process the effects of severe degradation.
In contrast, deep learning-based methods consistently demonstrate superior performance over traditional methods. On the Testing-E set, the method of Suh et al. [suh2022two] achieves both FM and p-FM values of 93.09 with a DRD of 3.12, while the method of Ju et al. [ju2024three] further improves these results to FM and p-FM values of 95.80 and 95.81, and reduces the DRD to 1.76. Our method achieves the best overall performance across all evaluation metrics, attaining FM and p-FM values of 98.11 and 98.14, a PSNR of 26.53 dB, a DRD of 0.82, and the highest ASM of 80.49.
When evaluated on the more challenging Testing-D set, deep learning-based methods maintain strong robustness. The FM score of the method of Suh et al. [suh2022two] decreases from 93.09 to 91.95, while the method of Ju et al. [ju2024three] exhibits a reduction from 95.80 to 94.86. In comparison, our method experiences only a slight decrease from 98.11 to 97.08, while still achieving the lowest DRD value of 1.38. Notably, compared with the method of Ju et al. [ju2024three], our method further reduces the DRD on Testing-D by approximately 39%, decreasing it from 2.27 to 1.38, while consistently outperforming competing methods in FM, p-FM, and PSNR.
Finally, in terms of model complexity, the methods of Suh et al. [suh2022two] and Ju et al. [ju2024three] contain 187.30M and 191.46M parameters, respectively, whereas our model requires only 31.22M parameters, achieving superior performance with a substantially reduced model size.